OpenStudy (anonymous):

A world record was set in a given year for the longest run on an ungaffed (fair) roulette wheel. The number 20 appeared 4 times in a row. What is the probability of the occurrence of this event? (Assume that there are 38 equally likely outcomes consisting of the number 1-36, 0, and 00. If you enter your answer in scientific notation, round the decimal value to two decimal places. Use equivalent rounding if you do not enter your answer in scientific notation.)

OpenStudy (anonymous):

help!

OpenStudy (aum):

There are 38 equally likely outcomes. The probability of number 20 appearing once is: \(\Large \frac{1}{38}\) The probability of number 20 appearing two times in a row is: \(\Large (\frac{1}{38})^2\) The probability of number 20 appearing three times in a row is: \(\Large (\frac{1}{38})^3\) The probability of number 20 appearing four times in a row is: \(\Large (\frac{1}{38})^4\) Use a calculator to evaluate \(\Large (\frac{1}{38})^4\). Express the result in scientific notation with the decimal value rounded to two decimal places.